OpenStudy (anonymous):

How do I set this up? I need to solve for n. 2 1/10 divided by "n" = 3 1/2

OpenStudy (anonymous):

2 1/10 =21/10 3 1/2=7/2 so 21n/10= 7/2 n=(7*10)/(2*21) =5/3 2 =1--- 3